OpenStudy (anonymous):

Please Help!!! Simplify showing each step. Tell what property you use in each step: 2(a + 4) + -8

OpenStudy (anonymous):

Any help is appreciated!!

hartnn (hartnn):

by distributive property a(b+c) = ab+ac so what will be 2(a+4) = ?

OpenStudy (anonymous):

2a + 8

hartnn (hartnn):

yes, + * - = - so, +-8 =- 8 now what is 2a+8-8 = ?

OpenStudy (anonymous):

2a

hartnn (hartnn):

thats your final answer.

OpenStudy (anonymous):

thanks!

hartnn (hartnn):

welcomes ^_^